OpenStudy (anonymous):

Please simply for medal: 6 7 --- - ---- v-2 2v+7

OpenStudy (nottim):

make the denominators the same first

OpenStudy (anonymous):

yea, but i don't know how to

OpenStudy (nottim):

whatever you do to the bottom, you do to the top.

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

what would the denomintor be?

OpenStudy (mertsj):

\[\frac{6}{y-2}-\frac{7}{2y+7}=\frac{6(2y+7)}{(y-2)(2y+7)}-\frac{7(y-2)}{(y-2)(2y+7)}\]

OpenStudy (nottim):

if you're asking a question at this level, you must know what im talking about.

OpenStudy (mertsj):

\[\frac{12y+42-7y+14}{(y-2)(2y+7)}=\frac{5y+56}{(y-2)(2y+7)}\]

OpenStudy (anonymous):

what level?

OpenStudy (nottim):

grade 9 math

OpenStudy (anonymous):

thankyou mertsj this is very helpful

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

where did you get the 5y+56?

OpenStudy (anonymous):

never mind i figuired it out

OpenStudy (mertsj):

I combined the like terms in the numerator. 12y-7y=5y 42+14=56

OpenStudy (anonymous):

thats the final answer right?